OpenStudy (anonymous):

n^2 - 4n - 1 = 0 Solve using the quadratic formula and simplify

OpenStudy (lgbasallote):

\[\large n = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[\Large \implies n = \frac{-(-4) \pm \sqrt{(-4)^2 - 4(1)(-1)}}{2(1)}\] does that help?

OpenStudy (anonymous):

Yes, thank you. I was struggling with these.